OpenStudy (anonymous):

A druggist has two mixtures of a certain chemical and water, one containing 10% of the chemical and the other containing 40% of the chemical. How much of each mixture should he use to make 10 ounces that contain 25% of the chemical?

OpenStudy (anonymous):

Really? We just answered this for you.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i need help please

OpenStudy (dumbcow):

x+y = 10 .1x + .4y = (.25)10

OpenStudy (anonymous):

7.5

OpenStudy (dumbcow):

?? what about 7.5 solve system using substitution y = 10-x --> .1x +.4(10-x) = 2.5 ... solve for x

OpenStudy (anonymous):

answer is 2.5

OpenStudy (dumbcow):

no...show your steps so i know what you did wrong

OpenStudy (anonymous):

okey 1

OpenStudy (anonymous):

@dumbcow my answer is 1

OpenStudy (dumbcow):

for x? no, by plugging 1 in you can see that both sides will not equal each other seems like you are guessing

OpenStudy (anonymous):

okey men isx ounces of 10% soln mixed with (10 – x) ounces of 40% soln. as total final volume is 10 ounces amount of chemical in x ounces = 0.1x amount of chemical in (10 – x) ounces = 0.4(10 – x) = 4 – 0.4x amount of chemical in 10 ounces = 10 × 0.25 = 2.5 chemical balance 0.1x + 4 – 0.4x = 2.5 0.3x = 1.5 x = 5 ounces of 10% mixed with (10 – 5) = 5 ounces of 40% to give 10 ounces of 25%

OpenStudy (dumbcow):

yes solution is (5,5) how do you go from guessing to writing a well laid out solution ...haha